OpenStudy (anonymous):

A right circular cone has a lateral area of 1,280 pi ft^2 and a base area of 512pi ft^2 . Find the volume.

OpenStudy (anonymous):

@amistre64 @SolomonZelman

OpenStudy (anonymous):

@phi

OpenStudy (phi):

do you know the formula for 1) lateral area 2) base area (area of a circle) 3) volume of a cone can you write down those 3 formulas?

OpenStudy (amistre64):

i always have to reinvent the lateral formula .... never sticks in my memory

OpenStudy (anonymous):

LA= pi*r*l V= 1/3*pi*r^2*h base= pi*r^2

OpenStudy (phi):

next, look at what we are given. base area of 512pi any ideas how to use that to find one of things we need to know to use the V formula?

OpenStudy (anonymous):

the radius?

OpenStudy (phi):

ok, sounds good. I would use \[ \pi r^2 = 512 \pi\] what do you get for r ?

OpenStudy (phi):

first divide both sides by pi

OpenStudy (anonymous):

sqroot of 512

OpenStudy (anonymous):

or 22.6

OpenStudy (phi):

I would leave things exact until the very end. so sqrt(512) you can write that as sqrt(2*256) and everybody knows 16*16 = 256 (kidding) so you can write r= 16 * sqr(2) or just keep it sqr(512)

OpenStudy (anonymous):

the radius again?

OpenStudy (phi):

if you know the radius (and you do) you put that on your list of "known" look at the formula: LA = pi * r * L which of those "letters" do we know? (btw , we know pi)

OpenStudy (anonymous):

la, pi, and r so it must be L?

OpenStudy (phi):

yes. You always try to get only 1 "unknown" letter. then you can "solve for it"

OpenStudy (phi):

can you solve for L (L is the slant height of the cone) ?

OpenStudy (phi):

do you need help solving for L ?

OpenStudy (anonymous):

I got 56.64

OpenStudy (anonymous):

h and L are different?

OpenStudy (phi):

let's see pi * r * L= 1280 r is sqr(512) \[ L= \frac{1280}{\pi \cdot \sqrt{512} } \] is that what you did to find L ?

OpenStudy (anonymous):

the pi will cancel

OpenStudy (phi):

oh, I did not see it was 1280 pi

OpenStudy (anonymous):

cuz the given is 1280 pi ft^2

OpenStudy (anonymous):

yup, 56.64 is what i got

OpenStudy (phi):

ok, I hope you are keeping track of these numbers. btw, if you use decimals, try to keep more decimals so you get a reasonably accurate answer.

OpenStudy (anonymous):

ok

OpenStudy (phi):

next, we look at the formula for volume of a cone V= 1/3*pi*r^2*h we know 1/3 the pi r^2 (notice this!) is the area of the base, and that is 512 pi we need h

OpenStudy (anonymous):

yup, how can we find h?

OpenStudy (phi):

do you see h, r and l form a right triangle ? any thoughts?

OpenStudy (anonymous):

i got 52 as the h

OpenStudy (phi):

how?

OpenStudy (anonymous):

h^2= L^2-r^2

OpenStudy (anonymous):

by PT

OpenStudy (phi):

ok, but keep more decimals

OpenStudy (anonymous):

i got a large number of it, i got 27798.96427

OpenStudy (anonymous):

is it right?

OpenStudy (phi):

write down the formula \[ V= \frac{1}{3} \pi r^2 h \] the pi r^2 is 512 pi (right?) h is 51.8459 so \[ V= \frac{1}{3} \pi \cdot 512 \cdot 51.8459 \] you need to multiply that out using a calculator.

OpenStudy (phi):

I get 27797.978 rounded to the nearest whole number 27,798

OpenStudy (anonymous):

yup i got a calculator here

OpenStudy (phi):

ok but you got 27798.96427 which rounds to 27799 (one too large?) typo?

OpenStudy (anonymous):

ok i will just use your answer

OpenStudy (phi):

it's not too hard to redo the multiplies, is it ?

OpenStudy (anonymous):

2 more problems here can you help me?

OpenStudy (phi):

maybe one more if you are quick?

OpenStudy (phi):

please make it a new post